# Elastic-plastic cell-based smoothed finite element method solving geotechnical problems

### v2

@@ 4 Construction of CSFEM elements
Table 1: Shape function value at different sites within an element (Fig. 2)
@@ 5 Implementation
5.1 Implementation of the elastic-plastic CSFEM
Fig. 3 and Algorithm 1 illustrate the computational workflow of the elastic–plastic analysis implemented in ABAQUS via the UEL interface. The main responsibility of the UEL subroutine is to evaluate and update the element contributions to the internal force vector (RHS) and stiffness matrix (AMATRX), based on the information supplied by the ABAQUS/Standard solver (Dassault Systèmes Simulia Corp., 2010). Within this framework, the stiffness matrix and the residual force vector are formulated as
@@ 5.1 Implementation of the elastic-plastic CSFEM
$\mathrm{RHS}=-\tilde{\mathrm{KU}}_{t+\Delta t}.$ (27)
@@
Figure 3: Framework sketch of implementing the elastic-plastic CSFEM within ABAQUS
Algorithm 1 Solving the CSFEM UEL
1:Node and element information, material properties, nodal displacement $u_{t}$
2:Nodal displacement $u_{t+1}$
@@
5:  solve nodal displacement $u^{k}_{t+1}$
@@
7:   solve shape functions $N_{k}(x)$ in Table 1
8:   solve strain–displacement matrix $\tilde{\mathrm{B}}_{c}$ in Eq.(18)
9:   solve elasto-plastic constitutive matrix $\mathrm{D}_{ep}$ in Eq.(25)
10:   solve stiffness matrix of subcells $\tilde{\mathrm{K}}_{c}$ in Eq.(16)
@@
12:  obtain element matrices $\tilde{\mathrm{K}}$
@@
16:solve nodal field variables $u_{t+1}=u^{k}_{t+1}$
5.2 Visualization of results
As shown in Fig. 4(a), the UEL in ABAQUS is represented by an “×” symbol, and its stress–strain results cannot be visualized directly. To address this limitation, a data-linking strategy was developed between the UEL and the user material subroutine (UMAT), as schematically illustrated in Fig. 3. Through this interface, the stress, strain, and other state variables (STATEV) computed within the UEL were transferred to the UMAT framework, allowing the use of standard ABAQUS elements for post-processing and visualization.
For visualization in ABAQUS/CAE, these quantities were exported via state variables declared in the input file using the *DEPVAR option and were stored at the integration points during the analysis. Consequently, as shown in Figs. 4(b) and (d), the numerical results obtained from the UEL could be accessed and visualized within the standard post-processing environment, providing clear field distributions comparable to those produced by conventional built-in elements.
Figure 4: Visualization of results
@@ 6.1 Pressurized thick-cylinder modeled using a quarter-annulus model
In this example, we considered a quarter-thickness cylindrical model subjected to internal pressure to verify the proposed method. The displacements in the polar coordinate system can be derived analytically (Timoshenko and Goodier, 1970).
@@
Figure 5: Fixed constraint beam mesh model at both ends
@@
Figure 7: Convergence of the relative error in the displacement.
Figure 8: Contour plots of results in thick-cylinder; (a) the FEM; (b) the CSFEM.
@@ 6.2 Biaxial test
For the present biaxial test, adopting the Mohr-Coulomb failure criterion with cohesion $c$ and friction angle $\varphi$, and taking the major and minor principal stresses as $\sigma_{1}$ and $\sigma_{3}$, respectively, the ultimate major principal stress is given by
$\sigma_{1,\mathrm{ult}}\;=\;\sigma_{3}\;\frac{1+\sin\varphi}{1-\sin\varphi}\;+\;\frac{2c\cos\varphi}{1-\sin\varphi}.$ (31)
@@ 6.3 Bearing capacity problem
In this example, we consider a flexible strip footing resting on the surface of a half space subjected to a uniform vertical load, as shown in Fig. 8(a). Due to symmetry, only half of the footing is modeled. Roller boundary conditions are imposed on the side edges, while the bottom boundary is fully fixed. The analysis is carried out under plane strain conditions using the mesh shown in Fig. 8(b), with 80 load increments of 8000 Pa. The corresponding soil properties are given in Table 2. For a soil with zero unit weight ($\gamma=0$), the ultimate bearing capacity of strip foundations can be expressed as:
@@
Table 2: Soil parameters of a flexible strip footing
@@
Table 3: Relative errors in bearing capacity between CSFEM and FEM
@@ 6.4 Tunnel excavation
A tunnel excavation problem is considered in this example, with the geometry shown in Fig. 15(a). The tunnel has a diameter of 2 m and is excavated in five stages, each 0.4 m thick. The model is constrained by fixed boundaries at the bottom and normal constraints on the lateral sides. The surrounding rock is characterized by an elastic modulus of $E=8$ MPa, Poisson’s ratio $\nu=0.27$, cohesion $c=120$ kPa, and an internal friction angle of $40^{\circ}$. The analysis begins with an in-situ stress equilibrium step, followed by excavation simulation using the element birth–death technique.
Fig. 16 shows the vertical displacement at the tunnel crown during the different excavation stages, from which it can be observed that CSFEM and FEM exhibit good agreement. When the excavation reaches the fourth stage, the vertical displacement nearly converges, with a maximum value of –0.021 m. Moreover, Fig. 17 and Fig. 18 present the vertical displacement and stress contour plots at the 1st, 3rd, and 5th excavation stages, further demonstrating the close agreement between CSFEM and FEM.
@@ 6.5 Slope stability analysis
To verify the accuracy of the proposed method, we examined a uniform slope with a height of $H=10\,\text{m}$, sloping at an angle of $\beta=45^{\circ}$ with a friction angle of $\phi=20^{\circ}$, a unit weight of $\gamma=20\,\text{kN/m}^{3}$, and a cohesion of $c=12.38\,\text{kPa}$, as shown in Fig. 19. Based on the soil properties, the slope has a factor of safety of exactly 1.0, as per the limit analysis solution presented by Chen (Chen and Liu, 2012). The side and bottom boundaries of slopes are constrained without displacement $(\Delta X=0;\,\Delta Y=0)$ and $(\Delta X=0;\,\Delta Y=0;\,\Delta Z=0)$, respectively.
Figure 19: Geometry and boundary conditions of the slope
The stability of slope is evaluated using the shear strength reduction technique (Matsui and San, 1992). The parameters of shear strength are defined as
$c_{m}=\frac{c}{F_{r}}$ (35)
$\varphi_{m}=\arctan\biggl(\frac{\tan\varphi}{F_{r}}\biggr)$ (36)
where $c$ and $\varphi$ are the actual shear strength parameters, $c_{m}$ and $\varphi_{m}$ are the shear strength which can maintain the stability of slope. $F_{r}$ is the strength reduction factor. In the reduction process, the input shear strength parameters $c$ and $\varphi$ are reduction to $c_{m}$ and $\varphi_{m}$, to trigger the slope failure. The failure criterion was defined by the big jump in the nodal displacement at a chosen point close to the slope’s surface.
We extracted the horizontal displacement of the slope vertex and plotted the relationship between horizontal displacement and the safety factor, as shown in Fig. 20. A big jump in displacement occurred in CS-FEM when the safety factor reached 1.02, whereas in FEM, the displacement jump occurred at a safety factor of 0.99. It is worth noting that the safety factor obtained by CS-FEM is very close to the value of 1.0 reported in the literature  (Dawson et al., 1999) for this slope. Moreover, Fig. 21 shows the distribution of plastic strain in the slope at failure. It can be seen from the figure that the plastic zones have fully penetrated when the slope fails.
Figure 20: The relationship between displacement and the FOS; (a) the CSFEM; (b) the FEM.
Figure 21: The distribution of plastic strain in the slope at failure ; (a) the CSFEM; (b) the FEM.
@@ 7 Conclusions
This paper presented an elastic–plastic cell-based smoothed finite element method (EP-CSFEM) tailored for geotechnical engineering. Based on the theoretical formulation, numerical implementation, and extensive verification, the following key conclusions can be drawn:
Enhanced accuracy and stability – The incorporation of strain smoothing significantly improves stress recovery and reduces mesh dependency, effectively mitigating volumetric locking in nearly incompressible soils.
Robust elastic–plastic performance – The proposed return-mapping scheme ensures accurate stress updates and stable convergence in nonlinear problems, even under large deformations.
Versatile applicability – Benchmark tests and practical examples (pressurized cylinder, biaxial soil test, strip footing, tunnel excavation, and slope stability) confirm that EP-CSFEM matches or surpasses conventional FEM in predicting displacements, stresses, and ultimate bearing capacities.
Implementation efficiency – The ABAQUS UEL framework enables direct output of stress–strain fields, facilitating seamless integration with standard post-processing tools.
Overall, the CSFEM provides a reliable and efficient numerical platform for analyzing complex soil–structure interaction and slope stability problems. Future work will extend the method to three-dimensional large-deformation analyses and coupled hydro-mechanical processes.

### v3

@@ 4 Construction of CSFEM elements
Table 1: Shape function value at different sites within an element (see Fig. 1).
@@ 5 Implementation
5.1 Implementation of the elasto-plastic CSFEM
Fig. 2 and Algorithm 1 illustrate the computational workflow of the elastic–plastic analysis implemented in ABAQUS via the UEL interface. The main responsibility of the UEL subroutine is to evaluate and update the element contributions to the internal force vector (RHS) and stiffness matrix (AMATRX), based on the information supplied by the ABAQUS/Standard solver (abaqus2010user). Within this framework, the stiffness matrix and the residual force vector are formulated as
@@ 5.1 Implementation of the elasto-plastic CSFEM
$\mathrm{RHS}=-\tilde{\mathrm{K}}\mathrm{U}_{t+\Delta t},$ (31)
@@
Figure 2: Framework of implementing the elasto-plastic CSFEM in ABAQUS using UEL and UMAT subroutines.
Algorithm 1 Solving the elasto-plastic CSFEM UEL
1:Node and element information, material properties, nodal displacement $\mathbf{u}_{t}$
2:Nodal displacement $\mathbf{u}_{t+1}$
@@
5:  solve nodal displacement $\mathbf{u}^{k}_{t+1}$
@@
7:   solve shape functions $N_{k}(x)$ in Tab.1
8:   solve strain–displacement matrix $\tilde{\mathbf{B}}_{c}$ in Eq.(18)
9:   solve elasto-plastic constitutive matrix $\mathbf{D}_{ep}$ in Eq.(24)
10:   solve stiffness matrix of subcells $\tilde{\mathbf{K}}_{c}$ in Eq.(16)
@@
12:  obtain element matrices $\tilde{\mathbf{K}}$
@@
16:solve nodal field variables $\mathbf{u}_{t+1}=\mathbf{u}^{k}_{t+1}$
5.2 UEL–UMAT data transfer mechanism for stress and strain visualization
As shown in Fig. 3, the UEL in ABAQUS is represented by an “×” symbol, and its stress–strain results cannot be visualized directly. To address this limitation, a data-linking strategy was developed between the UEL and the user material subroutine (UMAT), as schematically illustrated in Fig. 3. Through this interface, the stress, strain, and other state variables (STATEV) computed within the UEL were transferred to the UMAT framework, allowing the use of standard ABAQUS elements for post-processing and visualization.
For visualization in ABAQUS/CAE, these quantities were exported via state variables declared in the input file using the *DEPVAR option and were stored at the integration points during the analysis. Consequently, as shown in Fig. 3, the numerical results obtained from the UEL could be accessed and visualized within the standard post-processing environment, providing clear field distributions comparable to those produced by conventional built-in elements.
Figure 3: Data-linking scheme for transferring stress, strain, and other state variables from the UEL to UMAT-based visualization elements in ABAQUS.
To facilitate the transfer of stress, strain, and internal variables from the UEL to the UMAT framework, the state variables were explicitly declared in the input file through the *DEPVAR option. The corresponding ABAQUS syntax is listed below:
Listing 1: ABAQUS input for defining state variables via *DEPVAR
1*Material, name=Material-1
2*User Material, constants=0
3*Depvar
411,
51, EPSI, EPSI
62, TEPS, TEPS
73, PE11, PE11
84, PE22, PE22
95, PE33, PE33
106, PE12, PE12
117, S11, S11
128, S22, S22
139, S33, S33
1410, S12, S12
1511, PEMAG, PEMAG
The above definitions specify 11 state variables that store the quantities computed within the UEL at each integration point. For clarity, the physical meaning of these variables is summarized in Tab. 2.
Table 2: Definition of state variables used for UEL–UMAT data transfer.
Index Variable name Description
1 EPSI Equivalent plastic strain increment $\Delta\bar{\varepsilon}^{p}$
2 TEPS Total strain measure
3 PE11 Plastic strain component $\varepsilon^{p}_{11}$
4 PE22 Plastic strain component $\varepsilon^{p}_{22}$
5 PE33 Plastic strain component $\varepsilon^{p}_{33}$
6 PE12 Plastic strain component $\varepsilon^{p}_{12}$
7 S11 Stress component $\sigma_{11}$
8 S22 Stress component $\sigma_{22}$
9 S33 Stress component $\sigma_{33}$
10 S12 Stress component $\sigma_{12}$
11 PEMAG Total equivalent plastic strain $\bar{\varepsilon}^{p}$
@@ 6 Numerical examples
This section introduces several benchmark examples designed to verify the accuracy and convergence characteristics of the proposed elasto-plastic analysis framework. The numerical responses obtained with the CSFEM formulation are systematically compared with reference solutions generated using standard finite element analyses in ABAQUS. All computations were performed on a workstation equipped with an Intel Core i7-4710MQ processor (2.50 GHz) and 4 GB of RAM. The quality of the numerical results is quantified using the following relative error measure:
$\mathbf{e}_{L_{2}}=\frac{\left\|\mathbf{U}_{num}-\mathbf{U}_{ref}\right\|_{L_{2}}}{\parallel\mathbf{U}_{ref}\parallel_{L_{2}}},$ (32)
where $\mathbf{U}_{num}$ represents the numerical solution, and $\mathbf{U}_{ref}$ denotes the reference solution.
@@ 6.1 Pressurized thick-cylinder modeled using a quarter-annulus model
In this example, we considered a quarter-thickness cylindrical model subjected to internal pressure to verify the proposed method. The displacements in the polar coordinate system can be derived analytically (Timoshenko1970).
@@
Fig. 6 demonstrates that both the CSFEM and FEM exhibit satisfactory convergence behavior. As summarized in Tab. 3, the solution accuracy consistently improves with mesh refinement. In particular, for the finest mesh with a characteristic size of 0.003,m, the relative errors in the vertical displacement are $4.290\times 10^{-5}$ for the CSFEM and $1.144\times 10^{-4}$ for the FEM. Comparable trends are observed across all mesh levels, where the CSFEM yields smaller errors than the FEM. This indicates that the CSFEM offers enhanced accuracy relative to the conventional FEM. Furthermore, the displacement contour plots shown in Fig. 7 reveal an excellent agreement between the two methods, confirming the consistency of their numerical predictions.
Figure 4: Quarter model of a thick-walled cylinder subjected to internal pressure at the inner surface.
@@
(a) Relative error in $\mathrm{U_{x}}$
(b) Relative error in $\mathrm{U_{y}}$
Figure 6: Convergence of displacement errors in the (a) horizontal direction; (b) vertical direction.
Table 3: Relative errors in the displacement with mesh refining.
Size of element (m) Relative errors
Ux Uy
CSFEM FEM CSFEM FEM
0.125 $5.07\times 10^{-2}$ $5.186\times 10^{-2}$ $6.989\times 10^{-4}$ $1.860\times 10^{-3}$
0.0625 $2.650\times 10^{-2}$ $2.680\times 10^{-2}$ $1.817\times 10^{-4}$ $4.846\times 10^{-4}$
0.03 $1.270\times 10^{-2}$ $1.277\times 10^{-2}$ $4.290\times 10^{-5}$ $1.144\times 10^{-4}$
0.015 $6.340\times 10^{-3}$ $6.360\times 10^{-3}$ $1.051\times 10^{-5}$ $2.804\times 10^{-5}$
Figure 7: Contour plots of the numerical results for the thick cylinder problem: (a) results of the FEM; (b) results of the CSFEM.
@@ 6.2 Biaxial test
For the present biaxial test, adopting the Mohr-Coulomb failure criterion with cohesion $c$ and friction angle $\phi$, and taking the major and minor principal stresses as $\sigma_{1}$ and $\sigma_{3}$, respectively, the ultimate major principal stress is given by (FeiPeng2017AbaqusGeo)
$\sigma_{1,\mathrm{ult}}\;=\;\sigma_{3}\;\frac{1+\sin\phi}{1-\sin\phi}\;+\;\frac{2c\cos\phi}{1-\sin\phi}.$ (36)
@@
Figure 8: Schematic diagram for the biaxial test; (a) geometry and boundary conditions; (b) mesh model.
@@ 6.3 Bearing capacity problem
In this example, we consider a flexible strip footing resting on the surface of a half space subjected to a uniform vertical load, as shown in Fig. 11(a). Due to symmetry, only half of the footing is modeled. Roller boundary conditions are imposed on the side edges, while the bottom boundary is fully fixed. The analysis is carried out under plane strain conditions using the mesh shown in Fig. 11(b), with 80 load increments of 8000 Pa. The corresponding soil properties are given in Tab. 4. For a soil with zero unit weight ($\gamma=0$), the ultimate bearing capacity of strip foundations can be expressed as (sherif2012matlab):
@@
Table 4: Soil parameters of a flexible strip footing.
@@
Table 5: Relative errors in bearing capacity between CSFEM and FEM.
@@ 6.4 Tunnel excavation
A tunnel excavation problem is considered in this example, with the geometry shown in Fig. 14(a). The tunnel has a diameter of 2 m and is excavated in five stages, with each excavation step removing a thickness of 0.4 m. The model is constrained by fixed boundaries at the bottom and normal constraints on the lateral sides. The surrounding rock is characterized by an elastic modulus of $E=8$ MPa, Poisson’s ratio $\nu=0.27$, cohesion $c=120$ kPa, and an internal friction angle of $40^{\circ}$. The analysis begins with an in-situ stress equilibrium step, followed by excavation simulation using the element birth–death technique.
Fig. 15 shows the vertical displacement at the tunnel crown during the different excavation stages, from which it can be observed that CSFEM and FEM exhibit good agreement. When the excavation reaches the fourth stage, the vertical displacement nearly converges, with a maximum value of –0.021 m. Moreover, Fig. 16 and Fig. 17 present the vertical displacement and stress contour plots at the first, third, and fifth excavation stages, further demonstrating the close agreement between the CSFEM and FEM.
@@ 6.5 Slope stability analysis
To verify the accuracy of the proposed method, we examined a uniform slope with a height of $H=10\,\text{m}$, sloping at an angle of $\beta=45^{\circ}$ with a friction angle of $\phi=20^{\circ}$, a unit weight of $\gamma=20\,\text{kN/m}^{3}$, and a cohesion of $c=12.38\,\text{kPa}$, as shown in Fig. 18 (a). Based on the soil properties, the slope has a factor of safety of exactly 1.0, as per the limit analysis solution presented by Chen (chen2012limit). The side and bottom boundaries of slopes are constrained without displacement $(\Delta Y=0)$ and $(\Delta X=0;\,\Delta Y=0)$, respectively.
Figure 18: Schematic diagram for the slope stability analysis; (a) geometry and boundary conditions; (b) mesh model.
The stability of slope is evaluated using the shear strength reduction technique (matsui1992finite). The parameters of shear strength are defined as
$c_{m}=\frac{c}{F_{r}},$ (40)
$\phi_{m}=\arctan\biggl(\frac{\tan\phi}{F_{r}}\biggr),$ (41)
where $c$ and $\phi$ are the actual shear strength parameters, $c_{m}$ and $\phi_{m}$ are the shear strength which can maintain the stability of slope. $F_{r}$ is the strength reduction factor. In the reduction process, the input shear strength parameters $c$ and $\phi$ are reduction to $c_{m}$ and $\phi_{m}$, to trigger the slope failure. The failure criterion was defined by the big jump in the nodal displacement at a chosen point close to the slope’s surface.
We extracted the horizontal displacement of the slope vertex and plotted the relationship between horizontal displacement and the safety factor, as shown in Fig. 19. A big jump in displacement occurred in CSFEM when the safety factor reached 1.02, whereas in FEM, the displacement jump occurred at a safety factor of 0.99. It is worth noting that the safety factor obtained by CSFEM is very close to the value of 1.0 reported in the literature  (dawson1999slope) for this slope. Moreover, Fig. 20 shows the distribution of plastic strain in the slope at failure. It can be seen from the figure that the plastic zones have fully penetrated when the slope fails.
Figure 19: The relationship between displacement and the factor of safety (FOS); (a) the result of CSFEM; (b) the result of FEM.
Figure 20: The distribution of plastic strain in the slope at failure ; (a) the result of CSFEM; (b) the result of FEM.
@@ 7 Conclusions
An elasto-plastic cell-based smoothed finite element method (CSFEM) has been presented and implemented in ABAQUS for geotechnical analyses. The formulation replaces the compatible strain with a smoothed strain field evaluated over sub-cells and incorporates the Mohr-Coulomb model. The UEL-UMAT data-transfer mechanism is provided to allow direct visualization of stress, strain, and internal variables.
Three benchmark problems, one tunnel excavation, and one slope stability analysis were conducted to assess the performance of the proposed formulation. For the thick-cylinder, biaxial, and strip footing examples, the CSFEM consistently produces smaller displacement or load errors than the conventional FEM, confirming its higher accuracy in both elastic and elasto-plastic regimes. The tunnel excavation and slope stability examples further demonstrate the applicability of the method to staged construction and shear-strength reduction analyses, where the CSFEM and FEM provide comparable global responses, while the CSFEM still delivers well-resolved plastic strain patterns and failure mechanisms consistent with reference solutions.
Overall, the proposed CSFEM offers a simple and effective alternative to conventional FEM for elasto-plastic geotechnical problems. Future work will extend the method to three-dimensional analysis and more advanced constitutive models.